OpenStudy (anonymous):

I am lost on this word problem. Please help? (: A certain mix of concrete is composed of 94lb (one sack) of cement, 50 lb of water, 191lb of sand, and 299lb of gravel. The final mix of concrete will weight 151.2 lb/cu.ft. How many sacks of cement will be needed for a wall with a volume of 1760 cu.ft.

ganeshie8 (ganeshie8):

begin with writing down whats given

OpenStudy (anonymous):

Okay, I have drawn a visual as well.

ganeshie8 (ganeshie8):

good :) so the wall has a volume of1760 cu.ft you want to fill this volume with the concrete mix.

ganeshie8 (ganeshie8):

i need paper pen... wait :)

ganeshie8 (ganeshie8):

lets find out how much concrete is needed to fill 1760 cu.ft volume of wall

OpenStudy (anonymous):

okay

ganeshie8 (ganeshie8):

concrete needed for the entire wall = 1760 * 151.2 = ?

OpenStudy (anonymous):

=266112

ganeshie8 (ganeshie8):

concrete needed for wall = 26612 lb --------------------(1)

ganeshie8 (ganeshie8):

keep that aside for a sec

OpenStudy (anonymous):

Alright.

ganeshie8 (ganeshie8):

94lb (one sack) of cement, 50 lb of water, 191lb of sand, and 299lb of gravel.

ganeshie8 (ganeshie8):

add them all

OpenStudy (anonymous):

=634

ganeshie8 (ganeshie8):

94 + 50 + 191 + 299 = 634 lb of concrete mix

ganeshie8 (ganeshie8):

634 lb mix has 1 sack of cement

ganeshie8 (ganeshie8):

right ?

OpenStudy (anonymous):

I thought 94lb= 1 sack ?

ganeshie8 (ganeshie8):

thats correct ! 94lb of cement = 1 sack of cement

OpenStudy (anonymous):

So is every 94 pounds one sack ?

ganeshie8 (ganeshie8):

and in 634 lb of mixture, we have 94 lb of cement

ganeshie8 (ganeshie8):

we need to find ONLY cement sacks contained in the entire wall.

ganeshie8 (ganeshie8):

fore every 634 lb of mixture, 94 lb of cement is present

ganeshie8 (ganeshie8):

in 26612 lb of mixture, how much cement would be present ?

OpenStudy (anonymous):

Okay, I understand now. I thought: 634/94=6.7447 sacks

ganeshie8 (ganeshie8):

yeah good.... thats wrong as you see it now. that gives total sacks of mixture. but we are looking for cement only here

OpenStudy (anonymous):

Okay, I am not sure about the 26612 lb mixture.

ganeshie8 (ganeshie8):

so we have, 634 lb contains 1 sack of cement

ganeshie8 (ganeshie8):

is that clear upto this poitn ?

OpenStudy (anonymous):

Divide?

ganeshie8 (ganeshie8):

to find how many sacks of cement present in 26612 lb, you just divide 634 lb 26612 / 634 = ?

OpenStudy (anonymous):

=41.9748

OpenStudy (anonymous):

in sacks?

ganeshie8 (ganeshie8):

im getting 419.7 = ~ 420 sacks

ganeshie8 (ganeshie8):

oops! i was wrong... 41.97 = ~42 sacks is correct !!

OpenStudy (anonymous):

Okay

OpenStudy (anonymous):

Sorry, but where do I go from here? @ganeshie8

ganeshie8 (ganeshie8):

we are done lol. that wall requires ~42 sacks of cement.

OpenStudy (anonymous):

Oh okay haha. Thanks for your help!

ganeshie8 (ganeshie8):

np.... yw :)